OpenStudy (anonymous):

Need help The lengths of the sides of a triangle are, in units, 2-x over 2x+6,,, and 2x over 2x+6 and 1 over 2x+6 . What is the perimeter of the triangle in simplest terms (in units)? A. 1/2 B. 3+x over 2x+6 C.-1/2 D. 27+3x over 2x+6

OpenStudy (xapproachesinfinity):

the perimeter is the sum of all side lengths you have all the sides just add them up

OpenStudy (xapproachesinfinity):

\(\Large P=\frac{2-x}{2x+6}+\frac{2x}{2x+6}+\frac{1}{2x+6}= \frac{2-x+2x+1}{2x+6}=\frac{x+3}{2x+6}\)